What is the discount yield, bond equivalent yield, and effective annual return on a $1 million T-bill that currently sells at 93 3/8 percent of its face value and is 70 days from maturity? (Use 360 days for discount yield and 365 days in a year for bond equivalent yield and effective annual return. Do not round intermediate calculations. Round your answers to 3 decimal places.
Answer #1

Current Price=1,000,000 * (93 + 3/8)% = 0.93375*$1,000,000 = $9,33,750

Discount yield (dy)= (Face value- Price)/Face Value)(360/No.of days from maturity)

= I (t bill, dy) =(($1000,000-$933,750)/$1000,000)(360/70) =34.071%

Bond Equivalent yield (bey) ,BEY = ($Par Value/$Price- 1)(365/days left to maturity).

= I (t bill, bey) = (($1000,000-$933,750)/$933,750)(365/70) = 36.995%

Effective annual return is the rate that when applied to the initial investment will give a future value equal to the value arrived at after the compounding process.

EAR = (1+ 0.36995/(365/70))^(365/70) -1 = 42.963%